1. There are 4 contestants in a bicycle race. In how many different orders can they cross the finish line?
a. 4 b. 10 c. 24 d. 16
2. Clayton has to read 3 books for English Class. He can choose from a list of 5 books. HOw many different combinations of 3 books are possible?
a. 30 b. 15 c. 10
3.Carly is making cupcakes with nuts.She can choose between 4 flavors of cupcakes,3 colors of frosting,& 2 kinds of nuts. She chooses one flavor, one color & one kind of nut. How many different cupcakes could be made? a.10 b.16 c.4 d.24
4.In a bag of candies there are 13 red candies,13 green candies,13 yellow candies,& 13 blue candies.If you choose 1 candy from the bag,what is the probability the candy will not be blue? a.3/4 b.2/3 c.1/4 d.1/2

Answer by stanbon(57387)

1. There are 4 contestants in a bicycle race. In how many different orders can they cross the finish line?
4! = 4*3*2*! = 24
a. 4 b. 10 c. 24 d. 16
-----------------------------------------
2. Clayton has to read 3 books for English Class. He can choose from a list of 5 books. HOw many different combinations of 3 books are possible?
5C3 = 5C2 = (5*4)/(1*2) = 10
a. 30 b. 15 c. 10
-----------------------------------------
3.Carly is making cupcakes with nuts.She can choose between 4 flavors of cupcakes,3 colors of frosting,& 2 kinds of nuts. She chooses one flavor, one color & one kind of nut. How many different cupcakes could be made?
# of cupcakes = 4*3*2 = 24
a.10 b.16 c.4 d.24
-----------------------------------------
4.In a bag of candies there are 13 red candies,13 green candies,13 yellow candies,& 13 blue candies.If you choose 1 candy from the bag,what is the probability the candy will not be blue?
P(not blue) = 26/39 = 2/3
a.3/4 b.2/3 c.1/4 d.1/2
======================================================
Cheers,
Stan H.